OpenStudy (anonymous):

You accidently throw your car keys horizontally at 8.0 m/s form a cliff 64m high how far is the keys from the kliff

OpenStudy (anonymous):

First, find how long they fell, t, from distance y = (1/2) g t^2. 64m = 0.5 (10m/s^2) t^2. t = sqrt[64/5) Next use t (in sec) to find horizontal distance x=v t, x=(6 m/s) t There you are.